### v3

@@ 4.4 Gravito-inertial asteroseismology: high-order g modes in moderate rotators of intermediate mass
Analysis methods based on the TAR with the aim to deduce the rotation frequency of Dor and SPB stars were developed by Van Reeth et al. (2015a) and Ouazzani et al. (2017). The latter authors compared predictions for from the asymptotic expression in Eq. (16), from numerical computations adopting the TAR, and from full 2D computations with the ACOR code. This led to the important conclusions that computed from the TAR for zonal and prograde modes differ only by a few percent from those obtained with ACOR, while retrograde modes may led to differences up to about 10%. The method to derive and based on the TAR was applied to a sample of 37 Dor stars by Van Reeth et al. (2016), with independent validation of the results by Christophe et al. (2018). A slightly different analysis procedure to derive was developed by Takata et al. (2020), with overall consistent outcomes for (Ouazzani et al. 2019), but larger diversity (although still consistent) for the somewhat less precise estimation. Pedersen et al. (2021) tackled the challenging modelling for a sample of 26 SPB stars, where mixing in the core boundary layer and in the envelope is of major importance.
@@
The asteroseismic procedure to deduce the internal rotation frequency of Dor and SPB stars as commonly applied now is illustrated for prograde dipole modes in Fig. 7. The higher the rotation, the steeper the slope in the versus diagrams for prograde modes. Indeed, such modes represent waves travelling along the rotation of the star, thus their periodic variation re-occurs faster in the line-of-sight for faster rotation. Hence the mode periods, and by implication the difference of mode periods for consecutive overtones, decrease. This creates the characteristic downward tilt in the bottom panels of Fig. 7 as predicted theoretically already by Bouabid et al. (2013) and put into practice by Van Reeth et al. (2015a, 2016); Ouazzani et al. (2017). Retrograde gravito-inertial modes have the opposite effect and lead to an upward trend characterised by a positive slope (see Van Reeth et al. 2015b, for numerous examples).
@@
The detected upward slopes of period spacing patterns in fast rotators laid the foundation for the discovery of toroidal Rossby modes in Dor stars by Van Reeth et al. (2016), aside from retrograde gravito-inertial modes. It is remarkable that Rossby modes were discovered in many Dor stars from Kepler data already in 2016, while they were only detected two years later in the Sun (Löptien et al. 2018) as discussed in Sect. 4.1.1. From their dedicated description of and search for Rossby modes, Saio et al. (2018) found them to occur in the fastest rotators among the Dor stars, as well as in pulsating Be stars with outburst phenomena (Huat et al. 2009; Neiner et al. 2012a; Balona et al. 2015; Rivinius et al. 2016). Although several of the Kepler SPB stars also reveal photometric outbursts as found by Van Beeck et al. (2021), these have smaller amplitudes than those of the pulsating Be stars and do not necessarily lead to H emission in their spectra as is the case for the Be stars. So far, Rossby modes have not yet been found in SPB stars but the outbursting SPB and pulsating Be stars do share the common property of non-linear mode couplings (Baade et al. 2016; Van Beeck et al. 2021). This detected behaviour and theoretical developments of non-linear asteroseismology as in Lee (2012, 2022); Van Beeck et al. (2023) offer the great future potential to also include mode amplitudes in forward seismic modelling, aside from the frequencies.
@@
A summary of the findings based on 4-yr Kepler light curves of 611 Dor stars in terms of stellar properties is available in Li et al. (2020b) and for the internal rotation properties in Fig. 6 of Aerts (2021, not repeated here). The evolutionary context of the g-mode pulsators summarised in that review paper revealed overwhelmingly that dwarfs of intermediate mass have nearly uniform rotation throughout their entire main-sequence phase (see also Van Reeth et al. 2018), irrespective of their mass and current rotation rate. Levels of differentiality for the rotation of Dor stars remain below about 10% (Li et al. 2020b), irrespective of the value of the rotation rate itself because the sample covers rates from zero to almost the critical value. The summary of the internal rotation frequencies provided in Aerts (2021) also coupled the angular momentum of intermediate-mass dwarfs to that of evolved low- and intermediate-mass (sub)giants. This revealed that all these stars transport and lose their angular momentum in such a way that the angular momentum of their helium-burning cores by the time they are red giants is in agreement with the those of subdwarfs and white dwarfs (Aerts et al. 2019).
@@
The capacity to estimate and from versus diagrams as illustrated in Fig. 7 along with the envelope rotation from p modes and/or surface rotation from rotational modulation (Van Reeth et al. 2018) laid the foundation of the new sub-field of gravito-inertial asteroseismology for moderate and fast rotators (Aerts et al. 2018). This was meanwhile applied to samples of F and B pulsators as a way to estimate their level and type of core boundary and envelope mixing. Calibrating the internal mixing is necessary to properly age-date such stars and to deduce the mass of the helium core they will produce by the time of their core hydrogen exhaustion. Gravito-inertial asteroseismology nowadays delivers typical precisions between 5% and 20% for the age and helium core mass at the TAMS (Mombarg et al. 2019, 2021; Pedersen et al. 2021; Pedersen 2022b, a; Fritzewski et al. 2023). It also provides overwhelming evidence of extra mixing in the near-core region of rotating stars, fully in line with conclusion drawn from eclipsing binaries (Tkachenko et al. 2020; Johnston 2021). It also showed that the envelope mixing is connected to the internal rotation but that rotation is not the only physical cause of the element transport (Pedersen et al. 2021; Pedersen 2022b; Mombarg et al. 2023).
@@
Numerically stable tools to compute the eigenvalues of Laplace’s tidal equations are available in Townsend (2020) and make the estimation of the near-core rotation frequency of Dor and SPB stars a high-precision observational science. Given that the internal rotation frequency is meanwhile also accessible from TESS CVZ data, period spacing patterns due to low-degree gravito-inertial modes are within reach for many more pulsators (Garcia et al. 2022b, a), including stars in young open clusters (Fritzewski et al. 2023, and Li et al., submitted). This opens up the territory of applying gravito-inertial asteroseismology to samples of open clusters with a variety of ages, rotation rates, metallicities, single and binary populations, and merger products with the near-future potential to calibrate the internal structure and angular momentum of young clusters at their birth (see Fig. 1).
@@
Aside from the near-core rotation frequency deduced from gravito-inertial modes in the radiative envelope of intermediate-mass stars, a small fraction of the Dor stars also offers a powerful observational signal to derive the rotation rate of the convective core itself. This was first achieved by Ouazzani et al. (2020) from a pertinent characteristic dip in the period spacing pattern of the rapidly rotating Dor pulsator KIC 5608334, pointing to mode coupling between a core inertial mode and an envelope gravito-inertial mode in the presence of fast rotation. Saio et al. (2021) subsequently hunted such inertial core modes and found the signal in the period spacing diagram for 16 of the 611 Dor stars in Li et al. (2020b). These observational findings are meanwhile readily understood in terms of theoretical expressions for a mode coupling coefficient between inertial core modes and gravito-inertial envelope modes for Dor stars derived by Tokuno & Takata (2022). Numerical computations of these coupling coefficients by Aerts & Mathis (2023) from the best characterised Dor stars are in agreement with the theory. Although it concerns only a small fraction of Dor stars, notably the most rapid rotators, the detection of this mode coupling is of major importance. Indeed, it opens up the fast rotating convective cores of F dwarfs for observational scrutiny and confirms that their rotation is almost the same everywhere in their interior, from the core across the entire radiative envelope. Aerts & Mathis (2023) showed that the coupling coefficients for SPB stars are smaller, explaining why such mode coupling is absent for these hotter pulsators.
@@
Another way to probe the rotating convective core was established for a few pulsating late-type Be stars whose modelling is based on their fast rotation and on the observed properties of their stochastically excited low-frequency gravito-inertial modes, without having identifications for the individual modes (Neiner et al. 2012b, a, 2020). These cases concern such rapid rotators that the Coriolis force cannot be treated perturbatively nor in the TAR. Full computations in 1D but treating the deformation up to second order in were therefore done for the Be pulsator HD 49330 by Neiner et al. (2020). The predictions of the oscillation modes were based on the formalism by Lee & Baraffe (1995), with the aim to explain the observed variability spectrum qualitatively, without fitting individual mode frequencies.
@@
All the observational findings on internal rotation of gravito-inertial mode pulsators give rise to asteroseismic and evolutionary calibrations of angular momentum from the stellar birth towards evolved phases. Such studies are ongoing for single field stars and involve the global effect of various rotationally induced processes (Pedersen 2022a; Mombarg et al. 2023) and/or core magnetic dynamo fields (Moyano et al. 2023). The latter are based on the magnetic Tayler instability developed by Fuller et al. (2019) as an improvement over the classical Tayler-Spruit dynamo (see also Takahashi & Langer 2021). Aerts et al. (2021, their Fig. 8) estimated convective and wave Rossby numbers, core boundary stiffness values, and internal magnetic field strengths for a sample of 63 well characterised gravito-inertial mode pulsators. Assuming a stable dipole field, they predict field strengths in the ranges G and G for SPB and Dor stars, respectively. For the Dor stars, this is fully in line with the predictions by Cantiello et al. (2016) and Stello et al. (2016) based on dipole mixed mode suppression in about 20% of the Kepler red giants. Direct detections of core magnetic fields from mode splittings or magnetic characteristics in dipole mode period spacing patterns as recently found in a small fraction of red giants (see Sect. 4.1.3) are not yet established for gravito-inertial mode pulsators. Yet, given the numerous recent theoretical formalisms developed to hunt for specific signatures (Prat et al. 2019, 2020; Van Beeck et al. 2020; Dhouib et al. 2022; Rui et al. 2023), it should only be a matter of time for true direct detections to be achieved. For now, an upper limit of order G was found for the rapidly rotating CoRoT SPB star HD 43317 from asteroseismic modelling by Buysschaert et al. (2018). Lecoanet et al. (2022) confirmed this result by confronting the mode suppression observed in the star with predictions from 3D MHD simulations.
@@
Aside from probing of internal magnetism in or near the core of intermediate-mass pulsators via their gravito-inertial modes, Rossby modes and rotational modulation detected in 162 so-called ‘hump and spike’ stars gives indirect evidence for magnetic fields in the envelope of these stars (Henriksen et al. 2023a, b). However, it is yet unclear whether the observed variability is due to magnetic spots on the stellar surface triggered by a dynamo field generated in sub-surface layers or is rather caused by overstable convective modes that excite resonant g modes with frequencies equal to the rotation rate of the convective core (Lee & Saio 2020; Lee 2021b, 2022). Further research on the interaction between time-dependent convection, rotation, and magnetism in the nonadiabatic outer envelope is required to fully exploit the detected variability of these fastest rotators.
@@
In conclusion, gravito-inertial asteroseismology of intermediate-mass stars already made internal rotation and will soon make internal magnetism of moderate and fast rotators an observational science. Exploitation of the detected and identified gravito-inertial modes is meanwhile applied to hundreds of dwarfs of intermediate mass in the galaxy, with large potential for asteroseismology of young open clusters less than a few Gyr (Fritzewski et al. 2023, and Li et al., submitted). Such applications of asteroseismology to clusters observed by TESS and PLATO will deliver tight constraints on the angular momentum at stellar birth for various birth masses and metallicity regimes in the not too distant future (cf. Fig. 1).
@@ 5 Adding accurate dynamical masses and radii to increase asteroseismic precision
Asteroseismic modelling of rotating pulsators is a high-dimensional regression problem to solve (Aerts et al. 2018), with the aim to get high-precision estimates of the free parameters occuring in the stellar models. Aside from the measured effective temperature and gravity from high-resolution spectroscopy, and the star’s luminosity from Gaia astrometry, it is beneficial to limit the parameter space from additional and preferably model-independent input parameters.
@@
Major advances on this aspect may come from (future) interferometric radii, but this is limited to relatively bright stars. Combined interferometry and asteroseismology was so far mainly applied to bright slowly rotating low-mass Sun-like pulsators (Cunha et al. 2007; Huber et al. 2012). Applications to fast rotators are scarce and could so far only rely on uninterrupted space photometric light curves of short duration, limiting identification of sufficient modes (cf. Rasalhague and Altair analysed by Monnier et al. 2010; Le Dizès et al. 2021, respectively). It is noteworthy that the g modes of the bright hybrid Sct/Dor pulsator Rasalhague discovered in its MOST data were already correctly interpreted as gravito-inertial modes by Monnier et al. (2010). In our definition, this pulsator is a moderate rotator with its (or ) for d. However, the low frequency resolution of the light curve prevented the detection of clear period spacing patterns and the derivation of . Moreover, the methodology described in Sect. 4.4 based on the observable diagnostic in Fig. 7 did not yet exist at that time. It is definitely worthwhile to revisit the few bright fast rotating pulsators with a good interferometric radius and to hunt for g modes in their space photometry. Moreover, the new SPICA instrument in commissioning at the CHARA array (Mourard et al. 2022) holds great promise to intertwine interferometry and asteroseismology for a large sample of pulsators in the not too distant future.
@@
A highly successful and already applied extra constraint on asteroseimic modelling comes from dynamical masses of detached eclipsing binaries. Binary stars are found in diverse configurations in terms of orbital characteristics and physical properties of their individual components. In this section we focus on the classes of astrometric and eclipsing binaries as ideal test beds to help asteroseismic modelling. The key point is that such objects allow us to infer the masses and radii of the individual components by only relying on the basic laws of binary dynamics. Hence, the masses and radii are model-independent and their precision, and more importantly accuracy, is limited solely by the quality of the observations.
@@
Torres et al. (2010) presented a sample of 95 binary systems (190 stars in total) in a detached configuration, that is, systems where both components are well within their respective Roche equipotential surfaces and do not interact with each other. Their masses and radii are measured to better than accuracy. This sample covers M and R. Serenelli et al. (2021) discussed state-of-the-art methods for accurate determinations of stellar masses and presented a list of 200 benchmark stars, among which binary systems, with relative mass accuracy between [0.3,2]% for the mass range M. For the sake of brevity, we do not consider every type of binary system in this review but restrict to cases with the most precise and accurate determinations of stellar masses and radii. A recent review on eclipsing binaries with pulsating components in the era of space-based mission is provided in Lampens (2021).
@@
Having an accurate mass measurement alone does not enable stellar modelling at high precision due to the multitude of free parameters and the complex correlation patterns involved in the process. However, when combined with asteroseismic modelling, binary observables can break some of the parameter degeneracies. Both components of the binary are coeval and were born with the same initial chemical composition. Moreover, high-precision inferences of the effective temperatures of the stars from spectroscopic methods are possible, thanks to the fact that the components’ surface gravities are known with high accuracy from the dynamical measurements of the stellar masses and radii.
@@
Positioning of the binary components in the HRD or Kiel diagram and demanding that both stars were born together from the same molecular cloud, one can assess if and how well stellar models are able to predict the dynamical masses and evolutionary paths of both stars. This way, binarity has been exploited in numerous studies in the literature, which demonstrate that current stellar models can rarely reproduce the accurately measured properties of individual binary components and that extra physical ingredients are needed to reconcile models and observables. This has led to the need of extra mixing, from relatively small-scale near-core mixing (often simply referred to as convective core overshooting, Schroder et al. 1997; Guinan et al. 2000; Lacy et al. 2008; Torres et al. 2014b; Tkachenko et al. 2014; Meng & Zhang 2014; Claret & Torres 2016, 2017, 2018, 2019; Higl & Weiss 2017; Pavlovski et al. 2022; Rosu et al. 2020, 2022) to large-scale mixing caused by rotation and/or internal gravity waves (e.g., Ryan & Deliyannis 1995; de Mink et al. 2009; Garland et al. 2017; Pavlovski et al. 2009; Pavlovski & Southworth 2009; Pavlovski et al. 2018, 2023). The need of extra near-core and/or envelope mixing comes from the need of more massive convective cores of the binary components (e.g., Tkachenko et al. 2020; Johnston 2021; Martinet et al. 2021; Rosu et al. 2022). Interactions between magnetic fields, convection, atomic diffusion, and stellar spots have also been invoked to bring the radii of the stars into agreement with model predictions (e.g., Birkby et al. 2012; Torres et al. 2014a; Higl & Weiss 2017; Cruz et al. 2018; Garrido et al. 2019; Morales et al. 2022).

### v4

@@ 4.4 Gravito-inertial asteroseismology: High-order g modes in moderate rotators of intermediate mass
Analysis methods based on the TAR with the aim of deducing the rotation frequency of Dor and SPB stars were developed by Van Reeth et al. (2015a) and Ouazzani et al. (2017). The latter authors compared predictions for from the asymptotic expression in Eq. (16), from numerical computations adopting the TAR, and from full 2D computations with the ACOR code. This led to the important conclusions that values computed from the TAR for zonal and prograde modes differ only by a few percent from those obtained with ACOR, while retrograde modes may led to differences up to about 10%. The method used to derive and based on the TAR was applied to a sample of 37 Dor stars by Van Reeth et al. (2016), with independent validation of the results by Christophe et al. (2018). A slightly different analysis procedure to obtain was developed by Takata et al. (2020), with overall consistent outcomes for (Ouazzani et al. 2019), but larger diversity (although still consistent) for the somewhat less precise estimation. Pedersen et al. (2021) tackled the challenging modelling for a sample of 26 SPB stars, where mixing in the core boundary layer and in the envelope is of major importance.
@@
The asteroseismic procedure used to deduce the internal rotation frequency of Dor and SPB stars as commonly applied now is illustrated for prograde dipole modes in Fig. 7. The higher the rotation, the steeper the slope in the versus diagrams for prograde modes. Such modes represent waves travelling along the rotation of the star, and thus their periodic variation re-occurs faster in the line of sight for faster rotation. Hence, the mode periods, and by implication the difference of mode periods for consecutive overtones, decrease. This creates the characteristic downward tilt in the bottom panels of Fig. 7, as predicted theoretically by Bouabid et al. (2013) and put into practice by Van Reeth et al. (2015a, 2016); Ouazzani et al. (2017). Retrograde gravito-inertial modes have the opposite effect and lead to an upward trend characterised by a positive slope (see Van Reeth et al. 2015b, for numerous examples).
@@
The detected upward slopes of period spacing patterns in fast rotators laid the foundation for the discovery of toroidal Rossby modes in Dor stars by Van Reeth et al. (2016), in additon to the detection of retrograde gravito-inertial modes. It is remarkable that Rossby modes were discovered in many Dor stars from Kepler data already in 2016, while they were only detected two years later in the Sun (Löptien et al. 2018), as discussed in Sect. 4.1.1. From their dedicated description of and search for Rossby modes, Saio et al. (2018) found them to occur in the fastest rotators among the Dor stars, and in pulsating Be stars with outburst phenomena (Huat et al. 2009; Neiner et al. 2012a; Balona et al. 2015; Rivinius et al. 2016). Although several of the Kepler SPB stars also reveal photometric outbursts, as found by Van Beeck et al. (2021), they have smaller amplitudes than those of the pulsating Be stars and do not necessarily lead to H emission in their spectra as is the case for the Be stars. Rossby modes have not yet been found in SPB stars, but the outbursting SPB and pulsating Be stars do share the common property of non-linear mode couplings (Baade et al. 2016; Van Beeck et al. 2021). This detected behaviour and theoretical developments of non-linear asteroseismology (e.g. Lee (2012, 2022); Van Beeck et al. (2024)) offer the great future potential to also include mode amplitudes in forward seismic modelling, aside from the frequencies.
@@
A summary of the findings based on 4 yr Kepler light curves of 611 Dor stars in terms of stellar properties is available in Li et al. (2020b) and for the internal rotation properties in Fig. 6 of Aerts (2021, not repeated here). The evolutionary context of the g-mode pulsators summarised in that review paper reveal overwhelmingly that dwarfs of intermediate mass have nearly uniform rotation throughout their entire main-sequence phase (see also Van Reeth et al. 2018), irrespective of their mass and current rotation rate. Levels of core-to-surface rotation ratios of Dor stars remain below about 10% (Li et al. 2020b), irrespective of the value of the rotation rate itself. The summary of the internal rotation frequencies provided in Aerts (2021) also coupled the angular momentum of intermediate-mass dwarfs to that of evolved low- and intermediate-mass (sub-)giants. This revealed that all these stars transport and lose their angular momentum in such a way that the angular momentum of their helium-burning cores, by the time they are red giants, is in agreement with those of sub-dwarfs and white dwarfs (Aerts et al. 2019).
@@
The ability to estimate and from versus diagrams, as illustrated in Fig. 7, along with the envelope rotation from p modes and/or surface rotation from rotational modulation (Van Reeth et al. 2018) laid the foundation for the new sub-field of gravito-inertial asteroseismology for moderate and fast rotators (Aerts et al. 2018). This was applied to samples of F and B pulsators as a way to estimate the level and type of core boundary and the envelope mixing. Calibrating the internal mixing is necessary to properly age-date such stars and to deduce the mass of the helium core they will produce by the time of their core hydrogen exhaustion. Gravito-inertial asteroseismology currently delivers a typical precision of between 5% and 20% for the age and helium core mass at the TAMS (Mombarg et al. 2019, 2021; Pedersen et al. 2021; Pedersen 2022b, a; Fritzewski et al. 2024b). It also provides overwhelming evidence of extra mixing in the near-core region of rotating stars, fully in line with the conclusion drawn from eclipsing binaries (Tkachenko et al. 2020; Johnston 2021). It also shows that the envelope mixing is connected to the internal rotation, but that rotation is not the only physical cause of the element transport (Pedersen et al. 2021; Pedersen 2022b; Mombarg et al. 2023).
@@
Numerically stable tools to compute the eigenvalues of Laplace’s tidal equations are available in Townsend (2020) and make the estimation of the near-core rotation frequency of Dor and SPB stars a high-precision observational science. Given that the internal rotation frequency is also accessible from TESS CVZ data, period spacing patterns due to low-degree gravito-inertial modes are within reach for many more pulsators (Garcia et al. 2022b, a), including stars in young open clusters (Fritzewski et al. 2024b; Li et al. 2024a). This opens up the territory of applying gravito-inertial asteroseismology to samples of open clusters with a variety of ages, rotation rates, metallicities, single and binary populations, and merger products with the near-future potential to calibrate the internal structure and angular momentum of young clusters at their birth (see Fig. 1).
@@
Aside from the near-core rotation frequency deduced from gravito-inertial modes in the radiative envelope of intermediate-mass stars, a small fraction of Dor stars also offer a powerful observational signal that can be used to derive the rotation rate of the convective core itself. This was first achieved by Ouazzani et al. (2020) from a pertinent characteristic dip in the period spacing pattern of the rapidly rotating Dor pulsator KIC 5608334, pointing to mode coupling between a core inertial mode and an envelope gravito-inertial mode in the presence of fast rotation. Saio et al. (2021) subsequently hunted for these inertial core modes, and found the signal in the period spacing diagram for 16 of the 611 Dor stars in Li et al. (2020b). These observational findings are readily understood in terms of theoretical expressions for a mode coupling coefficient between inertial core modes and gravito-inertial envelope modes for Dor stars derived by Tokuno & Takata (2022). Numerical computations of these coupling coefficients by Aerts & Mathis (2023) from the best characterised Dor stars are in agreement with the theory. Although it is pertinent to only a small fraction of Dor stars, notably the most rapid rotators, the detection of this mode coupling is of major importance. It opens up the fast-rotating convective cores of F dwarfs for observational scrutiny and confirms that their rotation is almost the same everywhere in their interior, from the core across the entire radiative envelope. Aerts & Mathis (2023) showed that the coupling coefficients for SPB stars are smaller, explaining why such mode coupling is absent for these hotter pulsators.
@@
Another way to probe the rotating convective core was established for a few pulsating late-type Be stars whose modelling is based on their fast rotation and on the observed properties of their stochastically excited low-frequency gravito-inertial modes, without having identifications for the individual modes (Neiner et al. 2012b, a, 2020). These cases are for stars that rotate so rapidly that the Coriolis force cannot be treated either perturbatively or in the TAR. Full pulsation computations in 1D, treating the rotational deformation up to second order in , were done for the Be pulsator HD 49330 by Neiner et al. (2020). The predictions of the oscillation modes were based on the formalism by Lee & Baraffe (1995), with the aim of explaining the observed variability spectrum qualitatively, without fitting individual mode frequencies.
@@
All the observational findings on the internal rotation of gravito-inertial mode pulsators give rise to asteroseismic and evolutionary calibrations of angular momentum from the stellar birth to the evolved phases. Such studies are ongoing for single field stars, and involve the global effect of various rotationally induced processes (Pedersen 2022a; Mombarg et al. 2023) and/or core magnetic dynamo fields (Moyano et al. 2023). The latter are based on the magnetic Tayler instability developed by Fuller et al. (2019) as an improvement on the classical Tayler-Spruit dynamo (see also Takahashi & Langer 2021). Aerts et al. (2021, their Fig. 8) estimated convective and wave Rossby numbers, core boundary stiffness values, and internal magnetic field strengths for a sample of 63 well-characterised gravito-inertial mode pulsators. Assuming a stable dipole field, they predict field strengths in the ranges G and G for SPB and Dor stars, respectively. For the Dor stars, this is fully in line with the predictions by Cantiello et al. (2016) and Stello et al. (2016) based on dipole mixed mode suppression in about 20% of the Kepler red giants. Direct detections of core magnetic fields from mode splittings or magnetic characteristics in dipole mode period spacing patterns, as recently found in a small fraction of red giants (see Sect. 4.1.3), have not yet been established for gravito-inertial mode pulsators. However, given the numerous recent theoretical formalisms developed to hunt for specific signatures (Prat et al. 2019, 2020; Van Beeck et al. 2020; Dhouib et al. 2022; Rui et al. 2024), it should only be a matter of time for true direct detections to be achieved. For now, an upper limit of the order of G was found for the rapidly rotating CoRoT SPB star HD 43317 from asteroseismic modelling by Buysschaert et al. (2018). Lecoanet et al. (2022) confirmed this result by confronting the mode suppression observed in the star with predictions from 3D MHD simulations.
@@
Aside from probing the internal magnetism in or near the core of intermediate-mass pulsators via their gravito-inertial modes, Rossby modes and rotational modulation detected in 162 stars known as ‘hump and spike’ stars gives indirect evidence of magnetic fields in the envelopes of these stars (Henriksen et al. 2023a, b). However, it is still unclear whether the observed variability is due to magnetic spots on the stellar surface triggered by a dynamo field generated in sub-surface layers or is rather caused by overstable convective modes that excite resonant g modes with frequencies equal to the rotation frequency of the convective core (Lee & Saio 2020; Lee 2021b, 2022). Further research on the interaction between time-dependent convection, rotation, and magnetism in the nonadiabatic outer envelope is required to fully exploit the detected variability of these fastest rotators.
@@
In conclusion, gravito-inertial asteroseismology of intermediate-mass stars already made internal rotation and will soon make the internal magnetism of moderate and fast rotators an observational science. Exploitation of the detected and identified gravito-inertial modes is applied to hundreds of dwarfs of intermediate mass in the galaxy, with a large potential for asteroseismology of young open clusters of less than a few billion years in age (Fritzewski et al. 2024b; Li et al. 2024a). Such applications of asteroseismology to clusters observed by TESS and PLATO will deliver tight constraints on the angular momentum at stellar birth for various birth masses and metallicity regimes in the not too distant future (see Fig. 1).
@@ 5 Adding accurate dynamical masses and radii to increase asteroseismic precision
Asteroseismic modelling of rotating pulsators is a high-dimensional regression problem to solve (Aerts et al. 2018); the aim is to obtain high-precision estimates of the free parameters occurring in the stellar models. Aside from the measured effective temperature and gravity from high-resolution spectroscopy, and the star’s luminosity from Gaia astrometry, it is beneficial to limit the parameter space from additional and preferably model-independent input parameters.
@@
Major advances on this aspect may come from (future) interferometric radii, but this is limited to relatively bright stars. To date, combined interferometry and asteroseismology has mainly been applied to bright slow-rotating low-mass Sun-like pulsators (Cunha et al. 2007; Huber et al. 2012). Applications to fast rotators are scarce and so far can only rely on uninterrupted space photometric light curves of short duration, limiting the identification of sufficient modes (see Rasalhague and Altair analysed by Monnier et al. 2010; Le Dizès et al. 2021; Rieutord et al. 2024, respectively). It is noteworthy that the g modes of the bright hybrid Sct–Dor pulsator Rasalhague discovered in its MOST data were already correctly interpreted as gravito-inertial modes by Monnier et al. (2010). In our definition, this pulsator is a moderate rotator with a value of (or ) for d-1. However, the low frequency resolution of the light curve prevented the detection of clear period spacing patterns and the derivation of . Moreover, the methodology described in Sect. 4.4 based on the observable diagnostic in Fig. 7 did not yet exist in 2010. It is definitely worthwhile to revisit the few bright fast-rotating pulsators with a good interferometric radius and to hunt for g modes in their space photometry. Moreover, the new SPICA instrument in development at the CHARA array (Mourard et al. 2022) holds great promise to intertwine interferometry and asteroseismology for a large sample of pulsators in the not-too-distant future.
@@
A highly successful and already applied extra constraint on asteroseimic modelling comes from the dynamical masses of detached eclipsing binaries. Binary stars are found in diverse configurations in terms of orbital characteristics and physical properties of their individual components. In this section we focus on the classes of astrometric and eclipsing binaries as ideal test beds to help asteroseismic modelling. The key point is that such objects allow us to infer the masses and radii of the individual components by only relying on the basic laws of binary dynamics. Hence, the masses and radii are model-independent and their precision, and more importantly accuracy, is limited solely by the quality of the observations.
@@
Torres et al. (2010) presented a sample of 95 binary systems (190 stars in total) in a detached configuration, that is, systems where both components are well within their respective Roche equipotential surfaces and do not interact with each other. Their masses and radii are measured to better than accuracy. This sample covers M⊙ and R⊙. Serenelli et al. (2021) discussed state-of-the-art methods for the accurate determination of stellar masses and presented a list of 200 benchmark stars, among which binary systems, with relative mass accuracy between [0.3,2]% for the mass range M⊙. For the sake of brevity, we do not consider every type of binary system in this review, but restrict our analysis to cases with the most precise and accurate determinations of stellar masses and radii. A recent review of eclipsing binaries with pulsating components in the era of space-based mission is provided in Lampens (2021).
@@
Having an accurate mass measurement alone does not enable stellar modelling at high precision due to the multitude of free parameters and the complex correlation patterns involved in the process. However, when combined with asteroseismic modelling, binary observables can break some of the parameter degeneracies. Both components of the binary are coeval and were born with the same initial chemical composition. Moreover, high-precision inferences of the effective temperatures of the stars from spectroscopic methods are possible, thanks to the fact that the components’ surface gravities are known to high accuracy from the dynamical measurements of the stellar masses and radii.
@@
By positioning the binary components in the HRD or Kiel diagram and demanding that the two stars were born together from the same molecular cloud, one can assess whether and how well stellar models are able to predict the dynamical masses and evolutionary paths of both stars. This way, binarity has been exploited in numerous studies in the literature; the studies demonstrate that current stellar models can rarely reproduce the accurately measured properties of individual binary components and that extra physical ingredients are needed to reconcile models and observables. This has led to the need for extra mixing, from relatively small-scale near-core mixing (often simply referred to as convective core overshooting, Schroder et al. 1997; Guinan et al. 2000; Lacy et al. 2008; Torres et al. 2014b; Tkachenko et al. 2014; Meng & Zhang 2014; Claret & Torres 2016, 2017, 2018, 2019; Higl & Weiss 2017; Pavlovski et al. 2022; Rosu et al. 2020, 2022) to large-scale mixing caused by rotation and/or internal gravity waves (e.g. Ryan & Deliyannis 1995; de Mink et al. 2009; Garland et al. 2017; Pavlovski et al. 2009; Pavlovski & Southworth 2009; Pavlovski et al. 2018, 2023). The need for extra near-core and/or envelope mixing comes from the need of more massive convective cores of the binary components (e.g. Tkachenko et al. 2020; Johnston 2021; Martinet et al. 2021; Rosu et al. 2022). Interactions between magnetic fields, convection, atomic diffusion, and stellar spots have also been invoked to bring the radii of the stars into agreement with model predictions (e.g. Birkby et al. 2012; Torres et al. 2014a; Higl & Weiss 2017; Cruz et al. 2018; Garrido et al. 2019; Morales et al. 2022).